### 17 Apr 2026 v1

@@ 3.5 Equal-time algebra and hierarchical source-kernel structure
the displayed equal-time obstructions disappear, the surviving constraint pair $(\phi_{1},\phi_{2})$ becomes first-class in the strict equal-time sense, and the standard machinery of gauge generators, Dirac brackets, and reduced dynamics applies unchanged. The remainder of the paper, Sec. 4, develops this reduced description in that restricted subcase and uses it to confront the Banerjee route with the gauge-invariant equation-level formulation of Ref. [12]. No claim about the unrestricted external-current theory is made in what follows; the sectorial caveat is declared once here and not repeated subsection by subsection.
@@ 4.3 Reduced Hamilton equations and Maxwell-like rewriting
Hence
@@
Equation (4.22) is not an independent constitutive law but a compact rewriting of the reduced Hamilton equation for $\pi^{s}$; the notation is organizational. The source sector $T^{s}$ already contains both gauge-invariant and explicitly gauge-dependent contributions, which anticipates the asymmetric on-shell comparison with the Euler–Lagrange equations: the Gauss sector is anchored directly in the secondary constraint, whereas the Ampère-like sector inherits the full source-dependent deformation of the reduced Hamiltonian evolution. The term $\eta_{00}(\boldsymbol{\pi}\cdot\boldsymbol{\theta})\boldsymbol{\pi}$ in $\mathbf{H}$ encodes the anisotropic deformation of the effective magnetic permeability induced by the NC structure: with $\eta_{00}=+1$ and $\boldsymbol{\pi}\to\mathbf{E}$ on shell, this becomes $(\mathbf{E}\cdot\boldsymbol{\theta})\mathbf{E}$, coupling the electric sector to the NC direction $\boldsymbol{\theta}$ already at the constitutive level. The commutative limit $\theta\to 0$ recovers $\mathbf{H}=\mathbf{B}$, as required.
This Maxwell-like representation should not be overinterpreted. It does not establish a full equivalence between the reduced canonical description and the unrestricted field equations with arbitrary external currents. Rather, it expresses the reduced Hamiltonian evolution of the surviving sector in a form that is convenient for comparison with the mapped Euler–Lagrange structure. Its meaning is therefore internal to the restricted sufficient first-class subcase.

### 20 Apr 2026 v2

@@ 3.5 Equal-time algebra and hierarchical source-kernel structure
the displayed equal-time obstructions disappear, the surviving constraint pair $(\phi_{1},\phi_{2})$ becomes first-class in the strict equal-time sense, and the standard machinery of gauge generators, Dirac brackets, and reduced dynamics applies unchanged.
To see what these conditions select, consider the concrete case of purely $x^{1}$$x^{2}$ non-commutativity, $\theta^{12}=-\theta^{21}=\theta$ with all other components zero. The kernel $\Theta^{l}=\theta^{li}\partial_{i}J^{0}$ then requires $\partial_{1}J^{0}=\partial_{2}J^{0}=0$: the charge density must be uniform across the non-commutative plane. Likewise, $\Sigma^{lk}=0$ forces all spatial current components to be independent of $x^{1}$ and $x^{2}$, and $\Xi^{l}=0$ demands the same of $\partial_{\mu}J^{\mu}$. Any source of the form
$J^{\mu}=\bigl(\rho(t,x^{3}),\,0,\,0,\,j(t,x^{3})\bigr),\qquad\partial_{0}\rho+\partial_{3}j=0,$ (3.46)
automatically satisfies the first three conditions in Eq. (3.45). Under these conditions, the remaining requirement $\Phi_{3}^{\rm cand}\equiv 0$ reduces to ordinary current conservation $\partial_{\mu}J^{\mu}=0$, which is precisely the continuity equation already imposed in Eq. (3.46). The physical picture is therefore a conserved charge–current system whose profile may vary along the direction parallel to $\boldsymbol{\theta}$ but is uniform across the non-commutative plane. In particular, the sector conditions do not require $J^{\mu}$ to vanish: they restrict its derivative profile (spatial gradients and divergence) in the non-commutative directions, while the pointwise values $J^{0}$, $J^{k}$ remain non-trivial. This distinction is important for reading the reduced Hamilton equations of Sec. 4.3, where $J^{\mu}$ appears in its general functional form even though the brackets through which the equations are derived exist only for source profiles of the type (3.46).
The remainder of the paper, Sec. 4, develops this reduced description in that restricted subcase and uses it to confront the Banerjee route with the gauge-invariant equation-level formulation of Ref. [12]. No claim about the unrestricted external-current theory is made in what follows; the sectorial caveat is declared once here and not repeated subsection by subsection.
@@ 4.3 Reduced Hamilton equations and Maxwell-like rewriting
A remark on the logical status of the source terms is in order. The Dirac bracket $\{,\}_{*}$ constructed in Sec. 4.2 exists only in the restricted sector (3.45), because the invertibility of the constraint matrix in its standard Coulomb-gauge form requires $\Theta^{l}=\Sigma^{lk}=\Xi^{l}=0$. The canonical Hamiltonian $H_{c}=H_{\text{ws}}+H_{J}$, however, retains the full source coupling: the sector conditions constrain the admissible profiles of $J^{\mu}$—its spatial derivatives and divergence—but do not set $J^{\mu}$ itself to zero. The resulting Hamilton equations therefore contain the external current in its general functional form, while remaining valid only for source profiles satisfying Eq. (3.45). For concreteness, the admissible class includes the conserved, NC-plane-uniform profiles illustrated in Eq. (3.46): the source terms $J^{0}$, $J^{k}$ in the equations below are non-vanishing but have restricted spatial gradients. The equations are presented in this unsimplified form to make the comparison with the Euler–Lagrange system of Sec. 4.4 as direct as possible; no further reduction using the sector conditions on the source terms is attempted here.
@@
Two structural features of the source sector should be noted at the outset. First, the source terms in Eqs. (4.14) and (4.15) contain the gauge potential $A_{j}$ explicitly—not only through the field strength $F_{jl}$. In $\Lambda^{c}$, the term $\tfrac{1}{2}\eta_{00}J^{0}(\mathbf{A}\times\boldsymbol{\theta})^{c}$ depends on $A_{j}$ itself; in $\dot{\pi}^{s}$, the combination $J^{j}\theta^{sl}\partial_{j}A_{l}$ involves $\partial_{j}A_{l}$ rather than $F_{jl}=\partial_{j}A_{l}-\partial_{l}A_{j}$. These terms are the Hamiltonian counterpart of the gauge-non-covariant source coupling identified at the Lagrangian level by Adorno et al. [12]. Within the present gauge-fixed formulation ($A_{0}=0$, $\partial_{k}A^{k}=0$), the $A_{j}$ that appears is the physical transverse potential and the equations are self-consistent; but the appearance of the undifferentiated potential—rather than the field strength—in the source terms signals that the underlying action-level realization is not gauge covariant when external currents are present. The precise decomposition of this non-covariance into a scheme-dependent gauge-invariant piece and an explicitly gauge-dependent piece is carried out in the on-shell comparison of Sec. 4.5. Second, the field-strength bilinear sector (all terms not involving $J^{\mu}$) is manifestly gauge invariant, as it depends only on $F_{ij}$, $B^{i}$, and $\pi^{i}$. The asymmetry between gauge-invariant kinematics and gauge-non-covariant source coupling will be reflected in the on-shell comparison, where the Gauss sector is recovered exactly while the Ampère sector retains irreducible residues.
@@
Note that the last term, $\tfrac{1}{2}\theta^{ji}A_{j}J^{0}$, makes the effective electric field explicitly dependent on the gauge potential in the presence of sources; in the source-free limit $J^{0}\to 0$ this term vanishes and $E^{i}$ becomes a function of $F_{\mu\nu}$ and $\pi^{k}$ alone. Hence
@@
Equation (4.22) is not an independent constitutive law but a compact rewriting of the reduced Hamilton equation for $\pi^{s}$; the notation is organizational. The source sector $T^{s}$ contains the external current in its general functional form; the sector conditions (3.45) constrain only the derivative profile of $J^{\mu}$, not its pointwise values, so $J^{0}$, $J^{j}$, $J^{s}$ are not simplified here. Even so, $T^{s}$ already contains both gauge-invariant and explicitly gauge-dependent contributions, which anticipates the asymmetric on-shell comparison with the Euler–Lagrange equations: the Gauss sector is anchored directly in the secondary constraint, whereas the Ampère-like sector inherits the full source-dependent deformation of the reduced Hamiltonian evolution. The term $\eta_{00}(\boldsymbol{\pi}\cdot\boldsymbol{\theta})\boldsymbol{\pi}$ in $\mathbf{H}$ encodes the anisotropic deformation of the effective magnetic permeability induced by the NC structure: with $\eta_{00}=+1$ and $\boldsymbol{\pi}\to\mathbf{E}$ on shell, this becomes $(\mathbf{E}\cdot\boldsymbol{\theta})\mathbf{E}$, coupling the electric sector to the NC direction $\boldsymbol{\theta}$ already at the constitutive level. The commutative limit $\theta\to 0$ recovers $\mathbf{H}=\mathbf{B}$, as required.
This Maxwell-like representation should not be overinterpreted. In particular, the source sector $T^{s}$ is not gauge covariant: it contains the gauge potential $A_{j}$ through the term $J^{j}\theta^{sl}\partial_{j}A_{l}$, which does not combine into $F_{jl}$. Accordingly, the Ampère-like equation (4.22) does not have the same gauge-invariance status as its commutative counterpart. The representation is nevertheless useful because it separates the gauge-invariant field-strength block ($\mathbf{H}$, $D^{s}$) from the gauge-non-covariant source block ($T^{s}$) in a form that makes the comparison with the mapped Euler–Lagrange structure (Sec. 4.4) and the on-shell gauge-covariance analysis (Sec. 4.5) as direct as possible. Its meaning is therefore internal to the restricted sufficient first-class subcase.